# Identifying treatment response subgroups in observational time-to-event data

### v1

@@ Abstract
Identifying patient subgroups with different treatment responses is an important task to inform medical recommendations, guidelines, and the design of future clinical trials. Existing approaches for subgroup analysis primarily focus on Randomised Controlled Trials (RCTs), in which treatment assignment is randomised. Furthermore, the patient cohort of an RCT is often constrained by cost, and is not representative of the heterogeneity of patients likely to receive treatment in real-world clinical practice. Therefore, when applied to observational studies, such approaches suffer from significant statistical biases because of the non-randomisation of treatment. Our work introduces a novel, outcome-guided method for identifying treatment response subgroups in observational studies. Our approach assigns each patient to a subgroup associated with two time-to-event distributions: one under treatment and one under control regime. It hence positions itself in between individualised and average treatment effect estimation. The assumptions of our model result in a simple correction of the statistical bias from treatment non-randomisation through inverse propensity weighting. In experiments, our approach significantly outperforms the current state-of-the-art method for outcome-guided subgroup analysis in both randomised and observational treatment regimes.
Data and Code Availability
Experiments are performed on a synthetic dataset and a real-world dataset from the Surveillance, Epidemiology, and End Results program [^1] Available at https://seer.cancer.gov/. For reproducibility, implementation of the method, synthetic dataset, and analyses are available on Github  [^2] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
@@ 1 Introduction
Understanding heterogeneous therapeutic responses between patient subgroups is the core of treatment guidelines and the development of new drugs. Identifying such subgroups is valuable to inform clinical trials by identifying subgroups not responding to existing drugs, and to direct healthcare resources to those who might benefit most, and away from those who may be most harmed [11]. For instance, subgroup analysis in the BARI trial of patients with coronary artery disease supported the use of coronary artery bypass graft over percutaneous interventions for patients with diabetes, and the opposite for patients without diabetes [18], shaping subsequent guidelines. Figure 1 illustrates this idea: two groups may present opposite treatment responses, and would benefit from different recommendations. Our work aims to uncover such subgroups.
Randomised controlled trials (RCTs) remain the gold standard for identifying subgroups of treatment effects. In RCTs, patients are randomly assigned to control or treated groups, allowing researchers to assess the impact of an intervention or treatment. However, RCTs are costly and time-consuming, with restricted patient cohorts unrepresentative of the real-world diversity of patients who may receive treatment and treatment strategies [16].
Our work introduces a novel approach that diverges from traditional RCT-based methodologies [34, 36] by leveraging routinely collected observational data to identify patient subgroups. Observational studies encompass larger and more diverse cohorts reflective of real-world practices, offering the potential to uncover subgroups of treatment responses, that could be missed in RCTs. Prior works [5, 9, 29] have leveraged observational data while addressing biases from non-random treatment assignments [4, 15, 14]. However, this body of literature primarily focuses on estimating (i) averaged treatment effects at the population level or (ii) individualised treatment effects, thereby overlooking the identification of treatment effect subgroups.
@@
Our work extends Neural Survival Clustering [19] to estimate treatment effects, introducing the novel Causal Neural Survival Clustering (CNSC). This methodology is the first neural network approach to simultaneously estimate treatment effects and discover subgroups in observational settings, addressing both censoring and non-random treatment assignments without parameterising the survival distribution or treatment effects.
Section 2 formalises the problem of treatment subgroup discovery and non-randomisation correction. Then, we introduce the proposed monotonic neural network implementation in Section 2. Finally, we evaluate the proposed methodology on two datasets in Section 3 and compare it with related approaches described in Section 4.
@@ 2 Method
In this section, we describe our method, referring to Appendix A for technical proofs.
@@ 2.1 Problem setup
Central to our problem is the additional variable $Z$ associated with the latent, unobserved subgroup membership. Following Jeanselme et al. [19], we aim to identify a pre-specified $K\in\mathbb{N}$ number of subgroups [^3] As a pre-specified number of subgroups may be a limitation in a real-world setting where we do not know the underlying grouping structure, we explore how to select this parameter based on the likelihood of the predicted outcomes in Section 3.. We assume that $Z$ depends on $X$ and influences $T^{\prime}$. As we are interested in recovering the survival functions associated with the latent subgroups, we ignore the potential dependence of $T^{\prime}$ on $X$, considering the individual survival distribution as a mixture of the different subgroups.
@@
One can maximise the log-likelihood associated with the observed realisation $\{x_{i},t_{i},d_{i}\}$ of the random variables $X,T,D$ for patient $i$ (we omit $i$ where redundant) to estimate the survival function (Eq. (1)). Under Assumption 1, Equation (2) describes this factual [^4] Referring to the likelihood of the observed (rather than counterfactual) realisations. log-likelihood.
@@
where $\theta$ is the set of parameters characterising the estimated survival function $S_{\theta}$. Here, patients with an observed event ($d_{i}=1$) contribute the negative derivative of the survival function to the likelihood (see [19] for derivation). Each censored patient ($d_{i}=0$) contributes the probability of not observing an event before $t_{i}$, i.e., $S_{\theta}(t_{i}\mid x_{i})$, under Assumption 1 of independence between the censoring time and the event of interest conditional on the covariates.
@@
Figure 2: Graphical representation between covariates ($X$) and outcomes ($T,D$). Realisations of dashed variables are unobserved, while $X$, $A$, $T$ and $D$ are observed. $Z$ stands for the latent subgroup structure explaining the event time under no treatment $T^{\prime}_{0}$ and under treatment $T^{\prime}_{1}$. $A$ is the assigned treatment. $C$ is the censoring time.
This assumption is realistic, as (i) clinicians base treatment recommendations on the patients’ covariates, and (ii) subgroups are unknown and would consequently not inform treatment. If subgroups were known, one would not need the proposed methodology to uncover novel subgroups.
@@
In RCTs, treatment $A$ and covariates $X$ are by design independent. Hence, the subset of the patients receiving either treatment regime is representative of the overall population. Relying on this observed subset to estimate the survival distribution under each regime is a valid estimate for the population. In other words, maximising the factual likelihood to estimate survival under both treatment regimes results in a valid estimate. However, covariates and treatment are not independent in observational studies in which treatment recommendations depend upon the observed covariates. Formally, $\mathbb{P}(A\mid X)\neq\mathbb{P}(A)$ in general. For example, clinicians might recommend more aggressive treatment to patients in more severe conditions. This absence of randomisation results in a covariate shift between the treated and non-treated populations [9] as their covariate distributions differ [6]. In this setting, estimating survival by maximising the factual likelihood is no-longer enough to estimate the counterfactual survival distribution.
Under Assumption 6, approaches such as re-weighting [48], or penalisation on the dissimilarity of learnt representations [22, 47], or a combination of these approaches [9, 13, 45] are remedies to estimate the counterfactual likelihood from the observed data by reducing the difference between the two treatment regimes’ populations when modelling the outcome of interest. Specifically, Shalit et al. [45] demonstrate that the negative log likelihood is upper-bounded by observable quantities. Here we extend their notations to the survival setting as
@@
where $l$ is the log-likelihood consisting of both factual and counterfactual log-likelihoods, $l_{F}^{*}$ is the factual log-likelihood (defined in Eq.(2)) weighted with an inverse propensity of treatment weighting for each patient, $q_{\Phi}^{A=a}=q(\Phi(X)\mid A=a)$ is the density function associated with the transformation $\Phi$ of the covariates $X$, the Integral Probability Metric (IPM) is a distance between distributions, and $\gamma$ is a positive constant.
@@
Note that, under the dependencies we assume, only the subgroup assignment ($\mathbb{P}(Z\mid X)$) depends on the covariates $X$ and could, consequently, play the role of the transformation $\Phi$. However, treatment is assumed independent of the group membership under Assumption 3, meaning that treatment rate does not differ between subgroups. This assumption results in the regularisation term being null, i.e. $\text{IPM}(q_{Z\mid X}^{A=0},q_{Z\mid X}^{A=1})=0$ (see App. A.2 for derivation).
As a consequence, our work focuses on the first term of the upper bound in Eq. (4) by re-weighting the factual likelihood with the patient’s propensity score [13, 45]. Using an estimator $\hat{p}_{A}(x):=\mathbb{P}(A=1\mid X=x)$ of the propensity score, we use a truncated propensity weighting [3] scheme to avoid unstable weights in which the weights $w_{i}$ are defined as
@@
where $i$ is the patient index, and $a_{i}$ is the realisation of $A$ for patient $i$. Using the factual likelihood from Eq. (2) and the weights $w_{i}$ defined in Eq. (5), we derive the upper-bound of the Negative Log-Likelihood (NLL), later used to train our model, as
@@
Figure 3: Causal Neural Survival Clustering architecture. Latent parameter $l_{k}$ characterising the subgroup $k$ is inputted in the monotonic network $M$ to estimate the cumulative hazard $\Lambda_{k}$ under both treatment regimes. $G$ assigns the probability to belong to each subgroup given the patient’s covariate(s) $x$. To tackle the challenge of treatment assignment bias, the network $W$ estimates the treatment propensity used to weigh the training likelihood. In this context, patient survival is the average of the weighted survival distributions across subgroups under the given treatment.
@@ 2.2 Estimating the quantities of interest with neural networks
Similar to [19], a multi-layer perceptron $G$ with final Softmax layer assigns each patient characterised by covariates $x$ its probability of belonging to each subgroup, characterised through a $K$-dimensional vector of probabilities.
@@
Each subgroup $k$ is represented by a vector $l_{k}\in\mathbbm{R}^{L}$ of dimension $L$, a latent parametrisation of the cumulative hazard functions. The vector $l_{k}$ is concatenated with $t$ and used as input to a neural network $M$ with monotonic positive outcomes [^5] To ensure this constraint, we apply a square function on all weights as proposed in [19, 20]., with a final SoftPlus layer to model the cumulative hazard functions under both treatment regimes $\Lambda_{k}(t):=(\Lambda_{k}(t\mid A=0),\Lambda_{k}(t\mid A=1))$. We use the following transformation to ensure that no probability is assigned to negative times, a limitation raised concerning previous monotonic neural networks [46]:
@@
By modelling these two cumulative intensity functions, we can estimate the Group Average Treatment Effect (GATE), $\hat{\tau}_{k}(t)$, in subgroup $k$ as
@@
Under treatment randomisation, as in RCTs, the previous components would accurately model the observed survival outcomes and identify subgroups of treatment effects by maximising the factual likelihood. As previously discussed, to account for the treatment non-randomisation in observational studies, we weigh the factual likelihood using the propensity score of a patient estimated through a multi-layer perceptron $W$ with a final sigmoid transformation as
@@ 2.3 Training procedure
First, the network $W$ is trained to predict the binary treatment assignment by minimising the cross-entropy of receiving treatment. Then, training all other components relies on minimising the weighted factual log-likelihood introduced in Eq. (6). The use of monotonic neural networks results in the efficient and exact computation of the log-likelihood as automatic differentiation of the monotonic neural networks’ outcomes provides the instantaneous hazards $\lambda_{k}$ necessary for computing the survival function derivative used in the log likelihood [19, 20, 40].
@@ 3 Experimental analysis
As counterfactuals are unknown in observational data, we adopt—as is common practice in this research field—a synthetic dataset in which underlying survival distributions and group structure are known. Appendix C accompanies these synthetic results with the analysis of heterogeneity in adjuvant radiotherapy responses for patients diagnosed with breast cancer. Our code for the synthetic dataset, the model, and the reproduction of all experimental results is available on Github [^6] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
@@ 3.1 Data generation
We generate a population of $30,000$ divided into $K=3$ subgroups. We draw 10 covariates from normal distributions and survival distributions for each treatment regime following Gompertz distributions [37] parameterised by group membership and individual covariates. This choice reflects a setting in which time-to-event distributions under treatment or control regimes are different functions of the covariates. This simulation is more likely to capture the complexity of real-world responses, contrary to traditional evaluations of subgroup analysis, which often assume a linear treatment response, i.e., a simple shift between treated and untreated distributions. Further, we implement two treatment assignment scenarios: a randomised assignment in which treatment is independent of patient covariates, similar to RCTs (Randomised), and one in which treatment is a function of the patient covariates as observed in observational studies (Observational). Finally, non-informative censoring times are drawn also following a Gompertz distribution. Details of the generative process of our synthetic dataset are deferred to App. B.1.

### v3

@@ Abstract
Identifying patient subgroups with different treatment responses is an important task to inform medical recommendations, guidelines, and the design of future clinical trials. Existing approaches for subgroup analysis primarily rely on Randomised Controlled Trials (RCTs), in which treatment assignment is randomised. RCTs’ patient cohorts are often constrained by cost, rendering them not representative of the heterogeneity of patients likely to receive treatment in real-world clinical practice. When applied to observational studies, subgroup analysis approaches suffer from significant statistical biases particularly because of the non-randomisation of treatment. Our work introduces a novel, outcome-guided method for identifying treatment response subgroups in observational studies. Our approach assigns each patient to a subgroup associated with two time-to-event distributions: one under treatment and one under control regime. It hence positions itself in between individualised and average treatment effect estimation. The assumptions of our model result in a simple correction of the statistical bias from treatment non-randomisation through inverse propensity weighting. In experiments, our approach significantly outperforms the current state-of-the-art method for outcome-guided subgroup analysis in both randomised and observational treatment regimes.
@@ 1 Introduction
Understanding heterogeneous therapeutic responses between patient subgroups is the core of treatment guidelines and the development of new drugs. Identifying such subgroups is valuable to inform clinical trials by identifying subgroups not responding to existing drugs, and to direct healthcare resources to those who might benefit most, and away from those who may be most harmed (Foster et al., 2011). For instance, subgroup analysis in the BARI trial of patients with coronary artery disease supported the use of coronary artery bypass graft over percutaneous interventions for patients with diabetes, and the opposite for patients without diabetes (Investigators, 1996), shaping subsequent guidelines. Figure 1 illustrates this idea: two groups may present opposite treatment responses, and would benefit from different recommendations. Our work aims to uncover such subgroups.
Randomised controlled trials (RCTs) remain the gold standard for identifying subgroups of treatment effects. In RCTs, patients are randomly assigned to control or treated groups, allowing researchers to assess the impact of an intervention or treatment. However, RCTs are costly and time-consuming, with restricted patient cohorts unrepresentative of the real-world diversity of patients who may receive treatment and treatment strategies (Hernán & Robins, 2016).
Our work introduces a novel approach that diverges from traditional RCT-based methodologies (Nagpal et al., 2022a; 2023) by leveraging routinely collected observational data to identify patient subgroups. Observational studies encompass larger and more diverse cohorts reflective of real-world practices, offering the potential to uncover subgroups of treatment responses, that could be missed in RCTs. Prior works (Bica et al., 2020; Curth et al., 2021; Louizos et al., 2017) have leveraged observational data while addressing biases from non-random treatment assignments (Benson & Hartz, 2000; Hernán & Robins, 2010; Hernán, 2018). However, this body of literature primarily focuses on estimating (i) averaged treatment effects at the population level or (ii) individualised treatment effects, thereby overlooking the identification of treatment effect subgroups.
@@
Our work introduces the first neural network approach to simultaneously estimate treatment effects and discover subgroups in observational settings, addressing both censoring and non-random treatment assignments without parameterising the survival distribution or treatment effects. Section 2 formalises the problem of treatment subgroup discovery and non-randomisation correction. Then, we introduce the proposed monotonic neural network implementation in Section 2. Finally, we evaluate the proposed methodology on two synthetic datasets in Section 3, an extensive set of alternative settings in App. B and a real-world example described in App. C and compare it with related approaches reviewed in Section 4.
@@ 2.1 Problem setup
Central to our problem is the additional variable $Z$ associated with the latent, unobserved subgroup membership. Following Jeanselme et al. (2022), we aim to identify a pre-specified $K\in\mathbb{N}$ number of subgroups [^1] As a pre-specified number of subgroups may be a limitation in a real-world setting where we do not know the underlying grouping structure, we explore how to select this parameter based on the likelihood of the predicted outcomes in Section 3.. We assume that $Z$ depends on $X$ and influences $T^{\prime}$. As we are interested in recovering the survival functions associated with the latent subgroups, we ignore the potential dependence of $T^{\prime}$ on $X$, considering the individual survival distribution as a mixture of the different subgroups.
@@
One can maximise the log-likelihood associated with the observed realisations $\{x_{i},t_{i},d_{i}\}$ of the random variables $X,T,D$ for patient $i\in[\![1,N]\!]$ with $N$ the total number of patients (we omit $i$ where redundant) to estimate the survival function (Eq. equation 1). Under Assumption 1, Equation equation 2 describes this factual [^2] Referring to the likelihood of the observed (rather than counterfactual) realisations. log-likelihood.
@@
where $\theta$ is the set of parameters characterising the estimated survival function $S_{\theta}$. Here, patients with an observed event ($d_{i}=1$) contribute the negative derivative of the survival function to the likelihood (see Jeanselme et al. (2022) for derivation). Each censored patient ($d_{i}=0$) contributes the probability of not observing an event before $t_{i}$, i.e., $S_{\theta}(t_{i}\mid x_{i})$, under assumption 1 of independence between the censoring time and the event of interest conditional on the covariates.
@@
This assumption relaxes the traditional assumption of treatment randomisation. Note that this assumption remains realistic, as (i) clinicians base treatment recommendations on the patients’ covariates, and (ii) if subgroups were known, one would not need the proposed methodology to uncover novel subgroups.
Figure 2: Graphical representation between covariates ($X$) and outcomes ($T,D$). Realisations of dashed variables are unobserved, while $X$, $A$, $T$ and $D$ are observed.
@@
In RCTs, treatment $A$ and covariates $X$ are by design independent. Hence, the subset of the patients receiving either treatment regime is representative of the overall population. Relying on this observed subset to estimate the survival distribution under each regime is a valid estimate for the population. In other words, maximising the factual likelihood to estimate survival under both treatment regimes results in a valid estimate. However, covariates and treatment are not independent in observational studies in which treatment recommendations depend upon the observed covariates. Formally, $\mathbb{P}(A\mid X)\neq\mathbb{P}(A)$ in general. For example, clinicians might recommend more aggressive treatment to patients in more severe conditions. This absence of randomisation results in a covariate shift between the treated and non-treated populations (Curth et al., 2021) as their covariate distributions differ (Bica et al., 2021). In this setting, estimating survival by maximising the factual likelihood is no-longer enough to estimate the counterfactual survival distribution.
Under assumption 6, approaches such as re-weighting (Shimodaira, 2000), or penalisation on the dissimilarity of learnt representations (Johansson et al., 2016; Shi et al., 2019), or a combination of these approaches (Curth et al., 2021; Hassanpour & Greiner, 2019; Shalit et al., 2017) are remedies to estimate the counterfactual likelihood from the observed data by reducing the difference between the two treatment regimes’ populations when modelling the outcome of interest. Specifically, Shalit et al. (2017) demonstrate that the negative log likelihood is upper-bounded by observable quantities. Here we extend their notations to the survival setting as
@@
where $l$ is the log-likelihood consisting of both factual and counterfactual log-likelihoods, $l_{F}^{*}$ is the factual log-likelihood (defined in equation 2) weighted with an inverse propensity of treatment weighting for each patient, $q_{\Phi}^{A=a}=q(\Phi(X)\mid A=a)$ is the density function associated with the transformation $\Phi$ of the covariates $X$, the Integral Probability Metric (IPM) is a distance between distributions, and $\gamma$ is a positive constant.
@@
Note that, under the dependencies we assume, only the subgroup assignment ($\mathbb{P}(Z\mid X)$) depends on the covariates $X$ and could, consequently, play the role of the transformation $\Phi$. However, treatment is assumed independent of the group membership under assumption 3, meaning that treatment rate does not differ between subgroups. This assumption results in the regularisation term being null, i.e. $\text{IPM}(q_{Z\mid X}^{A=0},q_{Z\mid X}^{A=1})=0$ (see App. A.2 for derivation). This assumption is consequently necessary to correctly estimate the subgroups’ treatment effect, without penalising the clustering structure.
As a consequence, our work focuses on the first term of the upper bound in equation 4 by re-weighting the factual likelihood with the patient’s propensity score (Hassanpour & Greiner, 2019; Shalit et al., 2017). Using an estimator $\hat{p}_{A}(x):=\mathbb{P}(A=1\mid X=x)$ of the propensity score, we use a truncated propensity weighting (Austin & Stuart, 2015) scheme to avoid unstable weights in which the weights $w_{i}$ are defined as
@@
where $i$ is the patient index, and $a_{i}$ is the realisation of $A$ for patient $i$. Using the factual likelihood from equation 2 and the weights $w_{i}$ defined in equation 5, we derive the upper-bound of the Negative Log-Likelihood (NLL), later used to train our model, as
@@
Figure 3: Causal Survival Clustering architecture. Latent parameter $l_{k}$ characterising the subgroup $k$ is inputted in the monotonic network $M$ to estimate the cumulative hazard $\Lambda_{k}$ under both treatment regimes. $G$ assigns the probability to belong to each subgroup given the patient’s covariate(s) $x$. To tackle the challenge of treatment assignment bias, the network $W$ estimates the treatment propensity used to weigh the training likelihood. In this context, patient survival is the average of the weighted survival distributions across subgroups under the given treatment.
@@ 2.2 Estimating the quantities of interest with neural networks
Similar to Jeanselme et al. (2022), a multi-layer perceptron $G$ with final Softmax layer assigns each patient characterised by covariates $x$ its probability of belonging to each subgroup, characterised through a $K$-dimensional vector of probabilities.
@@
Each subgroup $k$ is represented by a vector $l_{k}\in\mathbbm{R}^{L}$ of dimension $L$, a latent parametrisation of the cumulative hazard functions. The vector $l_{k}$ is concatenated with $t$ and used as input to a neural network $M$ with monotonic positive outcomes [^3] To ensure this constraint, we apply a square function on all weights as proposed in Jeanselme et al. (2022; 2023)., with a final SoftPlus layer to model the cumulative hazard functions under both treatment regimes $\Lambda_{k}(t):=(\Lambda_{k}(t\mid A=0),\Lambda_{k}(t\mid A=1))$. We use the following transformation to ensure that no probability is assigned to negative times, a limitation raised concerning previous monotonic neural networks (Shchur et al., 2020):
@@
By modelling these two cumulative intensity functions, we can estimate the Subgroup Average Treatment Effect, $\hat{\tau}_{k}(t)$, in subgroup $k$ as
@@
Under treatment randomisation, as in RCTs, the previous components would accurately model the observed survival outcomes and identify subgroups of treatment effects by maximising the factual likelihood. As previously discussed, to account for the treatment non-randomisation in observational studies, we weight the factual likelihood using the propensity score of a patient estimated through a multi-layer perceptron $W$ with a final sigmoid transformation as
@@ 2.3 Training procedure
First, the network $W$ is trained to predict the binary treatment assignment by minimising the cross-entropy of receiving treatment. Then, training all other components relies on minimising the weighted factual log-likelihood introduced in equation 6. The use of monotonic neural networks results in the efficient and exact computation of the log-likelihood as automatic differentiation of the monotonic neural networks’ outcomes readily provides the instantaneous hazards $\lambda_{k}$ necessary for computing the survival function derivative in the log likelihood (Jeanselme et al., 2022; 2023; Rindt et al., 2022).
@@ 3 Experimental analysis
As counterfactuals are unknown in observational data, we adopt—as is common practice in this research field—a synthetic dataset in which underlying survival distributions and group structure are known. Appendix C accompanies these synthetic results with the analysis of heterogeneity in real-world, adjuvant radiotherapy responses for patients diagnosed with breast cancer. Our code to produce the synthetic dataset, the model, and the reproduction of all experimental results is available on Github [^4] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
@@ 3.1 Data generation
We generate a population of $3,000$ equally divided into $K=3$ subgroups. We draw 10 covariates from normal distributions with different centres and survival distributions for each treatment regime following Gompertz distributions (Pollard & Valkovics, 1992) parameterised by group membership and individual covariates. Note that this setting relaxes the independence of outcomes and covariates assumed by the proposed model. This choice reflects a setting in which time-to-event distributions under treatment or control regimes are different functions of the covariates. This simulation is more likely to capture the complexity of real-world responses, in contrast to traditional evaluations of subgroup analysis which often assume a linear treatment response, i.e., a shift between treated and untreated distributions. Further, we implement two treatment assignment scenarios: a randomised assignment in which treatment is independent of patient covariates, similar to RCTs (Randomised), and one in which treatment is a function of the patient covariates as in observational studies (Observational). Finally, non-informative censoring times are drawn following a Gompertz distribution. Details of the generative process of our synthetic dataset are deferred to App. B.1.